# Entropy-Aware Structural Alignment for Zero-Shot Handwritten Chinese Character Recognition

### v1

@@ Abstract
Zero-shot Handwritten Chinese Character Recognition (HCCR) aims to recognize unseen characters by leveraging radical-based semantic compositions. However, existing approaches often treat characters as flat radical sequences, neglecting the hierarchical topology and the uneven information density of different components. To address these limitations, we propose an Entropy-Aware Structural Alignment Network that bridges the visual-semantic gap through information-theoretic modeling. First, we introduce an Information Entropy Prior to dynamically modulate positional embeddings via multiplicative interaction, acting as a saliency detector that prioritizes discriminative roots over ubiquitous components. Second, we construct a Dual-View Radical Tree to extract multi-granularity structural features, which are integrated via an adaptive Sigmoid-based gating network to encode both global layout and local spatial roles. Finally, a Top-K Semantic Feature Fusion mechanism is devised to augment the decoding process by utilizing the centroid of semantic neighbors, effectively rectifying visual ambiguities through feature-level consensus. Extensive experiments demonstrate that our method establishes new state-of-the-art performance, significantly outperforming existing CLIP-based baselines in the challenging zero-shot setting. Furthermore, the framework exhibits exceptional data efficiency, demonstrating rapid adaptability with minimal support samples.
@@
Chinese character recognition , Zero-shot learning , Information entropy , Radical semantic modeling , Structural alignment
@@ 1 Introduction
Handwritten Chinese Character Recognition (HCCR) has long been a fundamental challenge in the field of pattern recognition and document analysis. Unlike alphabetic scripts, Chinese characters are logographic, characterized by highly complex topological structures and a vast vocabulary size—the GB18030-2022 standard alone encodes over 80,000 characters [43]. While deep learning-based methods have achieved remarkable success on standard benchmarks with fixed characters [38, 33], they often struggle with the open-set recognition problem, where the model encounters characters during testing that were never seen during training. This limitation has spurred significant research interest in Zero-Shot Learning (ZSL) for HCCR, which aims to recognize unseen characters by transferring knowledge from seen classes via intermediate semantic representations [31, 3, 4].
@@
To address these challenges, we propose a novel framework titled Entropy-Aware Structural Alignment Network. Specifically, we first introduce an Entropy-Aware Position Embedding mechanism. By quantifying the information entropy of each radical based on its statistical frequency, we dynamically modulate the positional embeddings via multiplicative interaction, forcing the model to attend more to high-entropy, discriminative components while suppressing redundant high-frequency signals. Secondly, we upgrade the structural representation by constructing Dual-View Radical Trees, which explicitly aggregate multi-granularity structural features from two distinct perspectives: a parent-centric view capturing global layout dependencies and a child-centric view preserving local compositional details.
Furthermore, to achieve precise cross-modal alignment, we propose a Radical Semantic Matching Module. Unlike simple similarity metrics employed in prior work [3, 40], this module employs an adaptive Sigmoid-based GateFusion network to integrate heterogeneous structural features and utilizes a Cross-Modal Attention mechanism [29, 13] to perform deep reasoning between visual and semantic representations. Finally, we introduce a Top-K Semantic Feature Fusion strategy, which augments the decoder’s query with the centroid of semantic neighbors to improve robustness against handwriting ambiguity.
@@ 2.1 Zero-shot HCCR via Decomposition
Traditional HCCR methods predominantly rely on deep convolutional neural networks (CNNs) trained on large-scale datasets with fixed characters [45, 25, 33, 35]. While these fully supervised models achieve high accuracy on closed sets, they fail to generalize to the open-set nature of Chinese characters, where thousands of rare characters and newly coined glyphs exist outside standard training sets [43].
@@ 2.2 Radical Representation and Structure Modeling
Despite these structural advancements, a critical theoretical gap remains: the neglect of information entropy. Current structural encoders, typically based on Tree-LSTMs [28] or Graph Convolutional Networks (GCNs) [12], generally assume that all nodes in the character tree contribute equally to the final representation. From an information-theoretic standpoint [23], this is suboptimal. High-frequency radicals (e.g., structural particles like ‘口’) carry low information entropy and are less informative for distinguishing fine-grained characters, whereas low-frequency radicals often serve as key discriminators. To date, few studies have incorporated statistical entropy as a prior to dynamically modulate the attention of structural encoders, a gap our work aims to fill.
@@ 2.3 Cross-Modal Alignment and Matching Mechanisms
Recent advances in Vision-Language Pre-training (VLP) have demonstrated the power of Cross-Attention mechanisms for fine-grained image-text alignment. Models like CLIP [22] use contrastive learning for global alignment, while subsequent works like SCAN [13] and ALBEF [14] employ deep cross-attention to allow the model to dynamically query relevant semantic components based on visual cues. Cai et al. [1] introduced learnable tokens to bridge the gap between local radical features and global character embeddings. However, applying such deep interaction mechanisms to HCCR is non-trivial due to the hierarchical nature of Chinese characters. Simply concatenating visual and semantic features often leads to modality dominance, where the stronger modality (usually the pre-trained visual backbone) overshadows the weaker one [32]. Therefore, a sophisticated fusion strategy—one that can adaptively balance global structure, local detail, and entropy priors—is required. This motivates the design of our GateFusion and Multi-Stage Semantic Matching Module, which moves beyond simple similarity to perform explicit, robust reasoning.
@@ 3 Methodology
Specifically, the visual stream processes handwritten images through a Multi-grid 2D Elastic Deformation module to simulate realistic writing distortions, followed by a backbone network to extract visual feature maps. Simultaneously, the semantic stream encodes the IDS of character characters. Unlike conventional methods that treat radicals equally, we introduce an Entropy-Aware Position Embedding to weight radicals based on their information content and construct Dual-View Radical Trees to capture both global layout and local structural dependencies. Finally, the visual and semantic representations are aligned via a Radical Semantic Matching Module, which performs coarse-to-fine reasoning using GateFusion and Cross-Modal Attention.
@@ 3.1.2 Multi-grid 2D Elastic Deformation
Figure 2: Illustration of the proposed Multi-grid 2D Elastic Deformation. Unlike traditional 1D strip-based methods, our approach constructs a dense 2D elastic mesh over the radical image. As visualized, the control points $p_{m,n}$ (red dots) are independently perturbed in both horizontal and vertical directions based on Gaussian random sampling. This mechanism explicitly models realistic non-rigid variations, such as local stroke squeezing and perspective warping, thereby enriching the structural diversity of the training samples.
@@ 3.2 Entropy-Aware Position Embedding
To address this, we propose an Entropy-Aware Position Embedding (EAPE) mechanism. This module explicitly quantifies the information density of each radical and uses a multiplicative interaction mechanism to dynamically modulate the embedding space.
@@ 3.2.1 Information Entropy Quantification
Formally, let $\mathcal{V}=\{r_{1},r_{2},\dots,r_{M}\}$ be the vocabulary of all unique radicals in the dataset. We first compute the global occurrence frequency $N(r_{k})$ for each radical $r_{k}\in\mathcal{V}$ across the entire training corpus. The probability of occurrence $P(r_{k})$ is given by:
$P(r_{k})=\frac{N(r_{k})}{\sum_{j=1}^{M}N(r_{j})}$ (1)
Distinct from standard definitions that typically employ base-2 logarithms (bits), we adopt the natural logarithm to measure uncertainty (nats), aligning better with the optimization dynamics of the continuous representation space. The Information Entropy $H(r_{k})$ is defined as:
$H(r_{k})=-\ln\left(P(r_{k})\right)$ (2)
This formulation ensures that ubiquitous radicals yield a low entropy value, while rare, discriminative radicals yield a high entropy value.
Figure 3: Visualization of the Information Entropy Statistics for over 400 radicals in the vocabulary. Each rectangle represents a unique radical. The color spectrum indicates the information content: (1) Deep Red ($\text{Entropy}\approx 9$) corresponds to rare radicals with high information density; (2) Deep Blue ($\text{Entropy}\approx 2$) corresponds to frequently occurring radicals with low information content; (3) The dominance of Red hues ($\text{Entropy}\in[6,9]$) indicates that the majority of radicals possess significant discriminative power. This unbalanced distribution serves as the statistical basis for our Entropy-Aware Position Embedding.
@@ 3.2.2 Embedding Modulation and Aggregation
Let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be the semantic embedding of the $i$-th radical obtained from our fine-tuned CLIP encoder, and $\mathbf{p}_{i}\in\mathbb{R}^{d}$ be the standard learnable positional embedding. The entropy-aware representation $\mathbf{v}_{i}$ is computed as:
@@
where $\odot$ denotes the element-wise product (Hadamard product), and $H(r_{i})$ is the scalar entropy value derived in Equation 2. In this formulation, the positional signal is first scaled by the radical’s entropy intensity ($H(r_{i})$). High-entropy radicals result in stronger positional signals, while low-entropy radicals have their spatial influence dampened. This scaled position vector is then fused with the semantic embedding $\mathbf{e}_{i}$ via element-wise multiplication, effectively creating an attention mask that highlights discriminative structural components.
Finally, to capture the global structural signature of the character, we aggregate the sequence into a global representation vector $\mathbf{v}_{ent}$ by averaging the entropy-weighted embeddings across the sequence length $L$:
$\mathbf{v}_{ent}=\frac{1}{L}\sum_{i=1}^{L}\mathbf{v}_{i}$ (4)
This mean-pooling operation ensures that the final representation is robust to sequence length variations while preserving the centroid of the entropy-weighted semantic space. This aligns the network’s focus with human cognitive intuition, ensuring that unique structural details contribute more significantly to the final character representation.
@@ 3.3 Dual-View Structural Representation
Figure 4: Illustration of the Dual-View Radical Tree Modeling using the character “深” (deep). (Left) Parent-centric Global View: This view models the bottom-up information aggregation. The root node ‘
@@ 3.3.4 Multi-Source Feature Encoding
    Entropy-Aware Representation ($\mathbf{V}_{ent}$): The global entropy-weighted feature derived in Equation 4, capturing the saliency of discriminative radicals.
@@ 3.4 Radical Semantic Matching Module
Having obtained the deformation-robust visual features from the CNN backbone and the dual-view structural embeddings from the radical tree, the final and most critical challenge is to align these two heterogeneous modalities. Preliminary works typically rely on shallow metric learning, such as calculating the Cosine similarity between the global visual vector and the radical embedding.
However, such linear alignment mechanisms are insufficient for zero-shot HCCR. They fail to capture the complex, non-linear correspondence between specific visual stroke patterns (e.g., a ”hook” stroke) and abstract radical semantics. To address this, we propose a Multi-Stage Semantic Matching Module, which explicitly performs coarse-to-fine reasoning through adaptive feature fusion, cross-modal attention, and neighbor-aware aggregation.
@@ 3.4.1 Adaptive GateFusion Network
Figure 5: Architecture of the Adaptive GateFusion Network. The module takes four heterogeneous structural embeddings ($\mathbf{V}_{ent},\mathbf{F}_{depth},\mathbf{F}_{parent},\mathbf{F}_{child}$) as input. Each branch undergoes a linear alignment followed by a Sigmoid-based gating mechanism to determine its element-wise contribution. The gated features are aggregated via summation, and the radical content embedding ($\mathbf{F}_{code}$) is finally injected to preserve the fundamental semantic identity. The output $\mathbf{P}_{sem}$ represents the fused semantic prototypes used for character recognition. This design allows the model to dynamically prioritize critical structural views while maintaining robust radical recognition.
Let $\mathcal{S}$ denote the input feature set, comprising the Entropy-Aware Position Embedding ($\mathbf{V}_{ent}$), the Depth Embedding ($\mathbf{F}_{depth}$), and the Dual-View Radical Tree features ($\mathbf{F}_{parent},\mathbf{F}_{child}$). Initially, to resolve the discrepancy in feature subspaces, each input feature $\mathbf{f}_{i}\in\mathcal{S}$ is projected into a unified high-dimensional semantic space $\mathbb{R}^{d}$ via a specific linear transformation $\tilde{\mathbf{f}}_{i}=\mathbf{W}_{i}\mathbf{f}_{i}+\mathbf{b}_{i}$. Subsequently, to achieve adaptive information flow, an independent gating unit employing the Sigmoid function $\sigma(\cdot)$ is applied to each aligned feature. This mechanism generates an element-wise importance map $\mathbf{g}_{i}=\sigma(\mathbf{W}_{gate,i}\tilde{\mathbf{f}}_{i}+\mathbf{b}_{gate,i})$, which modulates the feature magnitude through element-wise multiplication $\mathbf{h}_{i}=\mathbf{g}_{i}\odot\tilde{\mathbf{f}}_{i}$. By doing so, the network can autonomously emphasize discriminative structural cues while suppressing irrelevant noise.
@@ 4 Experiments
In this section, we conduct a comprehensive evaluation of the proposed Entropy-Aware Structural Alignment Network. We first detail the datasets and experimental protocols, followed by a comparative analysis against state-of-the-art (SOTA) methods on both zero-shot and few-shot recognition tasks. Furthermore, we perform extensive ablation studies to systematically isolate and verify the contributions of key components, specifically the entropy-aware position embedding, dual-view radical trees, and the multi-stage semantic matching module. Finally, qualitative visualizations are provided to offer deeper insights into the model’s internal reasoning mechanism and the learned feature space.
@@ 4.1 Experimental Setup
4.1.1 Datasets and Zero-shot Protocols
We evaluate our method using two standard benchmarks: CASIA-HWDB (1.0 and 1.1) [17] for training, and ICDAR 2013 [38] for zero-shot evaluation. These datasets contain offline handwritten samples produced by hundreds of writers, exhibiting significant variations in writing style, stroke thickness, and topology.
To strictly simulate the zero-shot learning scenario, we adopt the standard disjoint class partitioning protocol widely established in prior Zero-Shot HCCR works [3, 4, 11, 41, 19, 16, 21]. The experiments focus on the 3,755 most frequently used Chinese characters (GB2312 Level-1). We partition these categories into two disjoint sets ($\mathcal{Y}_{seen}\cap\mathcal{Y}_{unseen}=\emptyset$):
  • 1.
    Source Set (Seen): We select 2,755 character classes from the CASIA-HWDB 1.0 and 1.1 training sets. The model is trained exclusively on samples from these categories. Crucially, while the specific character combinations in the test set are withheld, this training set is chosen to ensure full coverage of the atomic radicals required to compose the unseen characters. This ”radical coverage” condition provides the physical basis for inferring unseen structures.
  • 2.
    Target Set (Unseen): We reserve the remaining 1,000 distinct character classes as the zero-shot testing set. During the inference phase, the model is evaluated on samples from the ICDAR 2013 competition dataset belonging to these unseen categories.
This setup challenges the model to recognize 1,000 entirely novel characters based solely on their learned structural definitions and visual-semantic alignment capabilities. Following standard evaluation metrics in the literature [43, 3], we primarily report the Top-1 Character Recognition Accuracy (Acc@1) to quantify performance.

### v2

@@ Abstract
Zero-shot Handwritten Chinese Character Recognition (HCCR) aims to recognize unseen characters by leveraging radical-based semantic compositions. However, existing approaches often treat characters as flat radical sequences, neglecting the hierarchical topology and the uneven information density of different components. To address these limitations, we propose an Entropy-Aware Structural Alignment Network that bridges the visual-semantic gap through information-theoretic modeling. First, we introduce an Information Entropy Prior to dynamically modulate positional embeddings via multiplicative interaction, acting as a saliency detector that prioritizes discriminative roots over ubiquitous components. Second, we construct a Dual-View Radical Tree to extract multi-granularity structural features, which are integrated via an adaptive Sigmoid-based gating network to encode both global layout and local spatial roles. Finally, a Top-K Semantic Feature Fusion mechanism is devised to augment the decoding process by utilizing the centroid of semantic neighbors, effectively rectifying visual ambiguities through feature-level consensus. Extensive experiments demonstrate that our method establishes new state-of-the-art performance, achieving an accuracy of 55.04% on the ICDAR 2013 dataset ($m=1500$), significantly outperforming existing CLIP-based baselines in the challenging zero-shot setting. Furthermore, the framework exhibits exceptional data efficiency, demonstrating rapid adaptability with minimal support samples, achieving 92.41% accuracy with only one support sample per class.
@@
Chinese character recognition , Zero-shot learning , Entropy-aware learning , Radical semantic modeling , Structural alignment
@@ 1 Introduction
Handwritten Chinese Character Recognition (HCCR) has long been a fundamental challenge in the field of pattern recognition and document analysis. Unlike alphabetic scripts, Chinese characters are logographic, characterized by highly complex topological structures and a vast vocabulary size—the GB18030-2022 standard alone encodes over 80,000 characters [42]. While deep learning-based methods have achieved remarkable success on standard benchmarks with fixed characters [37, 32], they often struggle with the open-set recognition problem, where the model encounters characters during testing that were never seen during training. This limitation has spurred significant research interest in Zero-Shot Learning (ZSL) for HCCR, which aims to recognize unseen characters by transferring knowledge from seen classes via intermediate semantic representations [30, 4].
@@
To address these challenges, we propose a novel framework titled Entropy-Aware Structural Alignment (EASA) network. Specifically, we first introduce an Entropy-Aware Position Embedding (EAPE) mechanism. By quantifying the information entropy of each radical based on its statistical frequency, we dynamically modulate the positional embeddings via multiplicative interaction, forcing the model to attend more to high-entropy, discriminative components while suppressing redundant high-frequency signals. Secondly, we upgrade the structural representation by constructing Dual-View Radical Trees, which explicitly aggregate multi-granularity structural features from two distinct perspectives: a parent-centric view capturing global layout dependencies and a child-centric view preserving local compositional details. Furthermore, to achieve precise cross-modal alignment, we propose a Radical Semantic Matching Module. Unlike simple similarity metrics employed in prior work [3, 39], this module employs an adaptive Sigmoid-based GateFusion network to integrate heterogeneous structural features and utilizes a Cross-Modal Attention mechanism [28, 12] to perform deep reasoning between visual and semantic representations. Finally, we introduce a Top-K Semantic Feature Fusion strategy, which augments the decoder’s query with the centroid of semantic neighbors to improve robustness against handwriting ambiguity.
@@ 2.1 Zero-shot HCCR via Decomposition
Traditional HCCR methods predominantly rely on deep convolutional neural networks (CNNs) trained on large-scale datasets with fixed characters [24, 34]. While these fully supervised models achieve high accuracy on closed sets, they fail to generalize to the open-set nature of Chinese characters, where thousands of rare characters and newly coined glyphs exist outside standard training sets [42].
@@ 2.2 Radical Representation and Structure Modeling
Despite these structural advancements, a critical theoretical gap remains: the neglect of information entropy. Current structural encoders, typically based on Tree-LSTMs [27] or Graph Convolutional Networks (GCNs) [11], generally assume that all nodes in the character tree contribute equally to the final representation. From an information-theoretic standpoint [22], this is suboptimal. High-frequency radicals (e.g., structural particles like ‘口’) carry low information entropy and are less informative for distinguishing fine-grained characters, whereas low-frequency radicals often serve as key discriminators.
@@ 2.3 Cross-Modal Alignment and Matching Mechanisms
Recent advances in Vision-Language Pre-training (VLP) have demonstrated the power of Cross-Attention mechanisms for fine-grained image-text alignment. Models like CLIP [21] use contrastive learning for global alignment, while subsequent works like SCAN [12] and ALBEF [13] employ deep cross-attention to allow the model to dynamically query relevant semantic components based on visual cues. Cai et al. [1] introduced learnable tokens to bridge the gap between local radical features and global character embeddings. However, applying such deep interaction mechanisms to HCCR is non-trivial due to the hierarchical nature of Chinese characters. Simply concatenating visual and semantic features often leads to modality dominance, where the stronger modality (usually the pre-trained visual backbone) overshadows the weaker one [31].
@@ 3.1.2 Multi-grid 2D Elastic Deformation
Figure 2: Illustration of the Multi-grid 2D Elastic Deformation. A dense 2D elastic mesh is constructed over the radical image, where control points $p_{m,n}$ (visualized as red dots) are independently perturbed.
@@ 3.2 Entropy-Aware Position Embedding
To address this, we propose an Entropy-Aware Position Embedding mechanism. This module explicitly quantifies the information density of each radical and uses a multiplicative interaction mechanism to dynamically modulate the embedding space.
@@ 3.2.1 Information Entropy Quantification
Formally, let $\mathcal{V}=\{r_{1},r_{2},\dots,r_{M}\}$ be the vocabulary of all unique radicals in the dataset. We first compute the global occurrence frequency $N(r_{k})$ for each radical $r_{k}\in\mathcal{V}$ across the entire training corpus. Specifically, the probability of the $k$-th radical is calculated as $P(r_{k})=N(r_{k})/\sum_{j=1}^{M}N(r_{j})$, where $N(r_{k})$ denotes the frequency of radical $r_{k}$ in the training corpus and $M$ is the vocabulary size. Distinct from standard definitions that typically employ base-2 logarithms (bits), we adopt the natural logarithm to measure uncertainty (nats), aligning better with the optimization dynamics of the continuous representation space. The Information Entropy is defined as $H(r_{k})=-\ln(P(r_{k}))$, which ensures that ubiquitous radicals yield a low entropy value while rare, discriminative radicals yield a high value.
Figure 3: Visualization of Information Entropy Statistics for over 400 radicals. The heatmap spectrum ranges from deep blue (low entropy) to deep red (high entropy), illustrating the non-uniform distribution of information density across the vocabulary.
@@ 3.2.2 Embedding Modulation and Aggregation
Let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ denote the semantic embedding of the $i$-th radical, retrieved from the pre-computed dictionary $\mathbf{E}_{rad}$ (defined in Section 3.1.3), and $\mathbf{p}_{i}\in\mathbb{R}^{d}$ be the standard learnable positional embedding. The entropy-aware representation $\mathbf{v}_{i}$ is computed as:
@@
where $\odot$ denotes the element-wise product (Hadamard product), and $H(r_{i})$ is the scalar entropy value. In this formulation, the positional signal is first scaled by the radical’s entropy intensity ($H(r_{i})$). High-entropy radicals result in stronger positional signals, while low-entropy radicals have their spatial influence dampened. This scaled position vector is then fused with the semantic embedding $\mathbf{e}_{i}$ via element-wise multiplication, effectively creating an attention mask that highlights discriminative structural components.
Finally, to capture the global structural signature, we aggregate the sequence into a global representation vector $\mathbf{v}_{ent}=(1/L)\sum_{i=1}^{L}\mathbf{v}_{i}$. This mean-pooling operation ensures robustness to sequence length variations while preserving the centroid of the entropy-weighted semantic space, aligning the network’s focus with human cognitive intuition.
@@ 3.3 Dual-View Structural Representation
Figure 4: Illustration of the Dual-View Radical Tree Modeling using the character “深” (deep). (Left) Parent-centric Global View: Depicts bottom-up aggregation, where the root ‘
@@ 3.3.4 Multi-Source Feature Encoding
    Entropy-Aware Representation ($\mathbf{V}_{ent}$): The global entropy-weighted feature captures the saliency of discriminative radicals.
@@ 3.4 Radical Semantic Matching Module
Having obtained the deformation-robust visual features from the CNN backbone and the dual-view structural embeddings from the radical tree, the final and most critical challenge is to align these two heterogeneous modalities. Preliminary works typically rely on shallow metric learning, such as calculating the Cosine similarity between the global visual vector and the radical embedding. However, such linear alignment mechanisms are insufficient for zero-shot HCCR. They fail to capture the complex, non-linear correspondence between specific visual stroke patterns (e.g., a ”hook” stroke) and abstract radical semantics. To address this, we propose a Multi-Stage Semantic Matching Module, which explicitly performs coarse-to-fine reasoning through adaptive feature fusion, cross-modal attention, and neighbor-aware aggregation.
@@ 3.4.1 Adaptive GateFusion Network
Figure 5: Architecture of the Adaptive GateFusion Network. The module aggregates four heterogeneous structural embeddings ($\mathbf{V}_{ent},\mathbf{F}_{depth},\mathbf{F}_{parent},\mathbf{F}_{child}$) via element-wise Sigmoid gating ($\sigma$) to modulate their importance. The gated features are summed, and the radical content embedding ($\mathbf{F}_{code}$) is explicitly injected as an additive semantic bias to preserve the fundamental identity, yielding the final prototype $\mathbf{P}_{sem}$.
Let $\mathcal{S}$ denote the input feature set, comprising the EAPE ($\mathbf{V}_{ent}$), the Depth Embedding ($\mathbf{F}_{depth}$), and the Dual-View Radical Tree features ($\mathbf{F}_{parent},\mathbf{F}_{child}$). Initially, to resolve the discrepancy in feature subspaces, each input feature $\mathbf{f}_{i}\in\mathcal{S}$ is projected into a unified high-dimensional semantic space $\mathbb{R}^{d}$ via a specific linear transformation $\tilde{\mathbf{f}}_{i}=\mathbf{W}_{i}\mathbf{f}_{i}+\mathbf{b}_{i}$. Subsequently, to achieve adaptive information flow, an independent gating unit employing the Sigmoid function $\sigma(\cdot)$ is applied to each aligned feature. This mechanism generates an element-wise importance map $\mathbf{g}_{i}=\sigma(\mathbf{W}_{gate,i}\tilde{\mathbf{f}}_{i}+\mathbf{b}_{gate,i})$, which modulates the feature magnitude through element-wise multiplication $\mathbf{h}_{i}=\mathbf{g}_{i}\odot\tilde{\mathbf{f}}_{i}$. By doing so, the network can autonomously emphasize discriminative structural cues while suppressing irrelevant noise.
@@ 4 Experiments
In this section, we conduct a comprehensive evaluation of the proposed Entropy-Aware Structural Alignment Network. We first detail the datasets and experimental protocols, followed by a comparative analysis against state-of-the-art (SOTA) methods on both zero-shot and few-shot recognition tasks. Furthermore, we perform extensive ablation studies to systematically isolate and verify the contributions of key components. Finally, qualitative visualizations are provided to offer deeper insights into the model’s internal reasoning mechanism.
@@ 4.1 Experimental Setup
4.1.1 Datasets and Protocols
We evaluate our method on the standard CASIA-HWDB (1.0/1.1) [16] and ICDAR 2013 [37] benchmarks. Following the strict zero-shot protocol established in prior works [10, 20], we focus on the 3,755 Level-1 GB2312 characters, partitioning them into two disjoint sets ($\mathcal{Y}_{seen}\cap\mathcal{Y}_{unseen}=\emptyset$). The Source Set (Seen) comprises 2,755 classes from CASIA-HWDB for training, ensuring full coverage of atomic radicals. The Target Set (Unseen) reserves the remaining 1,000 classes for zero-shot evaluation using samples from ICDAR 2013. We report Top-1 Character Recognition Accuracy (Acc@1) to quantify performance.